OpenStudy (anonymous):

simplify the expression x^-4y^3x^4y^2

OpenStudy (anonymous):

y^5, ( n^x * n^-x = 1 )

OpenStudy (anonymous):

x^-4y^3x^4y^2 = x^4 y^3 y^2/ x^4 = y^5